OpenStudy (anonymous):

The function f(x)=9/(1+49x^2) is represented as a power series Find the first few coefficients in the power series. I got the first few coefficients, but can't figure out the radius of convergence. Can someone help?

OpenStudy (anonymous):

\[f(x)=\frac{ 9 }{ (1+49^{2}) }\] gives \[9(1-(49x^2)+(49x^2)^2\] so c0= 9 c1= 0 c2= -441 c3= 0 c4= 21609 but how do i find the radius of convergence?

OpenStudy (anonymous):

@saiken2009

OpenStudy (anonymous):

i know my C's are right. i need the radius of convergence. thanks

OpenStudy (anonymous):

can someone please help? i didn't need help finding the C's. i already found them. the ones @UditKulka wrote are wrong anyway. please help

terenzreignz (terenzreignz):

@saiken2009 Hi ^_^ I don't know how to get the radius of convergence (normally) but when faced with a function of this form \[\Large \frac{a}{1\pm u}\] This is represented by the power series \[\Large a\sum_{n=0}^\infty(\mp u)^n\] BUT ONLY if \(\large |u|<1\) Now, it just so happens that in this case, your \(\large u = 49x^2\) can you work it out from here?